When steam at 100℃ is passed over a large block of ice , 64g of ice melts. What is the amount of steam condenses

Grade:11

1 Answers

Sahir
5 years ago
Let the amount of steam condense is M Then heat released in condensation= M×LL is latent heat of vaporisationL= 540 cal/ gm And this will be equal to the heat taken by the ice to melt 64 and latent heat of fusion for melting ice is 80 cal/ gmM×540= 64×80 M=9.48 gm